A titration is carried out between HCl (aq) and 0.276 M KOH(aq). There is 38.6 mL acid, and it takes 65.3 mL of the base to reach the endpoint of the titration. What is the concentration (M) of the acid ?
1
Expert's answer
2020-10-20T08:37:15-0400

Solution:

The reaction between hydrochloric acid (HCl) and potassium hydroxide (KOH) is:

HCl + KOH = KCl + H2O

According to the equation: n(HCl) = n(KOH)

Therefore,

CM(HCl) × V(HCl) = CM(KOH) × V(KOH)

CM(HCl) = CM(KOH) × V(KOH) / V(HCl)

CM(HCl) = (0.276 M × 65.3 mL) / (38.6 mL) = 0.4669 M = 0.467 M


Answer: The concentration of the acid (HCl) is 0.467 M.